### v2

# Enhancing Function-Calling Capabilities in LLMs: Strategies for Prompt Formats, Data Integration, and Multilingual Translation

Abstract
Large language models (LLMs) have significantly advanced autonomous agents, particularly in zero-shot tool usage, also known as function calling. This research delves into enhancing the function-calling capabilities of LLMs by exploring different approaches, including prompt formats for integrating function descriptions, blending function-calling and instruction-following data, introducing a novel Decision Token for conditional prompts, leveraging chain-of-thought reasoning, and overcoming multilingual challenges with a translation pipeline. Our key findings and contributions are as follows: (1) Instruction-following data improves both function-calling accuracy and relevance detection. (2) The use of the newly proposed Decision Token, combined with synthetic non-function-call data, enhances relevance detection. (3) A tailored translation pipeline effectively overcomes multilingual limitations, demonstrating significant improvements in Traditional Chinese. These insights highlight the potential for improved function-calling capabilities and multilingual applications in LLMs.
Enhancing Function-Calling Capabilities in LLMs: Strategies for
Prompt Formats, Data Integration, and Multilingual Translation
Yi-Chang Chen  Po-Chun Hsu  Chan-Jan Hsu  Da-shan Shiu MediaTek Research {yi-chang.chen, pochun.hsu, chan.hsu, ds.shiu}@mtkresearch.com
1 Introduction
The field of autonomous agents has seen remarkable advancements in recent years, largely driven by the capabilities of large language models (LLMs). These models have significantly enhanced the performance of autonomous agents across a variety of tasks Huang et al. (2024); Qin et al. (2024); Qu et al. (2024). A critical ability for these agents is zero-shot tool usage, also known as function calling. This capability allows LLMs to access up-to-date information from the internet or in-house databases and leverage third-party services, enabling integration with various systems. Such capabilities open up numerous potential applications, including electronic design automation Zhong et al. (2023), financial reporting Theuma and Shareghi (2024), and travel planning Hao et al. (2024).
Despite the progress made through tuning-based methods Grattafiori et al. (2024); Liu et al. (2024a, b) for enabling function-calling capabilities, there remains a gap in research regarding the format variance of prompts, the combination of function-calling data with instruction-following data, and multilingual limitations. This work aims to address these gaps by investigating the following aspects:
Prompt Formats: We explore two strategies for incorporating function descriptions into prompts: (1) introducing a dedicated role for presenting function descriptions, and (2) embedding function descriptions within the system role alongside usage instructions. We aim to determine the impact of these formats on function-calling performance.
Data Integration: We examine the combination of function-calling data with instruction-following data to assess its impact on both instruction-following and function-calling capabilities. Our findings indicate that the use of instruction-following data significantly enhances function-calling accuracy and relevance detection.
Decision Token: We propose a novel Decision Token for conditional prompts, designed to improve relevance detection and facilitate the creation of synthetic non-function-call data for fine-tuning. Our results show that the inclusion of the Decision Token and non-function-call data enhances function-calling relevance detection.
Chain-of-Thought (CoT) Reasoning: We incorporate CoT reasoning through a synthetic data pipeline that constructs reasoning descriptions from sequences of conversations and function calls.
Multilingual Translation: We address the multilingual limitations of current function-calling models by introducing a translation pipeline specifically tailored to overcome the challenges of direct translation methods. Our Traditional Chinese experiments confirm this approach’s effectiveness.
In summary, this research provides valuable insights into enhancing LLMs’ function-calling capabilities and highlights the potential for multilingual applications. The following sections detail our methodology, experiments, and results, demonstrating the effectiveness of our proposed strategies.
Figure 1: An illustration of prompt templates used for function calling and instruction following in LLMs. During training, LLMs are given conditional prompts (shown on the left) and tasked with generating corresponding text completions (shown on the right). When a function call is required, the model generates structured function calls in the form of a list of functions, where each function is specified with its arguments in the format func_name(arg1=value1, ...). Special tokens, including <s>, <|im_start|>, <|im_end|>, <|answer|>, and <|use_tool|>, are each represented by a single token after tokenization. For more details, refer to Section 3.1.
2 Related Work
Integrating function-calling capabilities into LLMs significantly broadens their problem-solving abilities by enabling interactions with external tools and APIs. Studies have shown that API-integrated LLMs can perform tasks such as programming assistance Gao et al. (2022), real-time information retrieval Schick et al. (2023), complex mathematical computations He-Yueya et al. (2023), and internet utilization Komeili et al. (2021); Gur et al. (2024). This allows LLMs to access up-to-date information and leverage third-party services, facilitating integration with various systems across advanced applications like electronic design automation Zhong et al. (2023), financial reporting Theuma and Shareghi (2024), and travel planning Hao et al. (2024).
To enable such function-calling capabilities, researchers have explored two main categories of methods. The first involves sophisticated prompting techniques. Frameworks like ReACT Yao et al. (2022) and its successors Xu et al. (2023); Shinn et al. (2023); Yang et al. (2023b); Crouse et al. (2024); Wang et al. (2024) combine reasoning and acting within prompts to guide model responses.
More closely related to our work, the second category focuses on training models to generate function calls through fine-tuning. Fine-tuned models such as Gorilla Patil et al. (2023), ToolAlpaca Tang et al. (2023), ToolLlama Qin et al. (2024), and the Hermes 3 series by Nous-Research Teknium et al. (2024) enhance function-calling capabilities by relying on synthetic data generated by proprietary models like GPT-4 or ChatGPT. Open-source initiatives like NexusRaven-V2 Nexusflow.ai (2023) and IBM’s Granite-20B-FunctionCalling Abdelaziz et al. (2024) aim to develop function-calling models suitable for commercial use without relying on proprietary data. Moreover, many works involve self-supervision to further enhance performance across diverse domains Schick et al. (2023); Parisi et al. (2022); Yang et al. (2023a); Liu et al. (2024a).
Among the works in the second category, some fine-tuned models and datasets have been openly released. For instance, ToolAlpaca Tang et al. (2023) and ToolLLM Qin et al. (2024) have made available their synthetic data or data generation pipelines. ToolACE Liu et al. (2024a) has released both the fine-tuned Llama model and the self-instruction dataset. Additionally, the Gorilla team developed a comprehensive benchmark to evaluate LLMs’ function-calling capabilities Yan et al. (2024).
Notably, ToolACE Liu et al. (2024a) demonstrated that diversified function-calling sample data helps models learn better function-calling abilities. However, there is a lack of comprehensive analysis on how variations in prompt and meta-information design, as well as the impact of non-function-calling-related instruction tuning data, affect the effectiveness of function-calling capabilities. Existing studies tend to adopt specific prompt templates without extensively investigating the impact of different designs, indicating a need for further research in this area.
3 Methodology
3.1 Prompt Templates for Function Calling and Instruction Following
We employ a tuning-based approach to enable both function-calling and instruction-following capabilities in our LLMs. This involves fine-tuning pretrained base models using prompt templates based on the Chat Markup Language (ChatML), a widely adopted format introduced by OpenAI.
Two main strategies for incorporating function descriptions into prompts are explored: (1) introducing a dedicated role, such as tools, to represent function descriptions in JSON format (Figure 1(b)); and (2) embedding function descriptions alongside usage instructions within the system role (Figure 1(c)). In the latter strategy, both instruction-following and function-calling are guided by the system prompt.
During training, the LLMs are provided with conditional prompts as described above and are tasked with generating appropriate text completions. Based on the context, the fine-tuned model dynamically decides whether to respond directly or invoke functions. If no relevant functions are available, the model directly answers the query (Figure 1(d)). Otherwise, if function calls are needed, the model generates structured function calls in the form of a list of functions (Figure 1(f)).
In the experiments, we investigated the performance comparison of different conditional prompts and the use of training data across various metrics for instruction-following and function-calling capabilities, as discussed in Section 4.2.
3.2 Decision Token
Achieving high performance in relevance detection is challenging, often hindered by the scarcity of negative samples in most synthetic datasets Liu et al. (2024a, b).
To address this, we propose the novel Decision Token mechanism. LLMs generate responses through next-token prediction, where each step involves a classification task to select the next token. The Decision Token concept leverages the fact that each token prediction is essentially a classification. By introducing a pair of special tokens, the model can predict a binary classification that determines whether to answer the query directly or invoke function calls before generating a detailed response or function calls, respectively. Specifically, this process introduces a pair of special tokens, <|answer|> and <|use_tool|>, as shown in Figure 1(e) and (g). If the model chooses to provide a direct answer, it outputs <|answer|> first; if it chooses function calling, it outputs <|use_tool|> first. This classification task forces the model to make a decision based on the user query and provided functions before delving into the details, thereby enhancing the stability of its output.
The Decision Token also facilitates the creation of non-function-call data from function-called data. To generate non-function-call data, consider an example where the original data involves three functions: func_A, func_B, and func_C. Based on the user query, func_A is helpful and thus called in the original data point. By assuming that func_B and func_C are not helpful, we can create non-function-call data by removing func_A as input. With only func_B and func_C as the remaining functions, function calling should not be triggered from user query and a direct answer should be provided. This allows us to easily obtain non-function-call data. Previously, generating non-function-call data for training was challenging because it required specific LLM responses for non-function-call cases. However, with the Decision Token, we can train the model to output only <|answer|> in non-function-call cases. During inference, this is not an issue because the model will continue to provide an appropriate response after <|answer|>.
The experiments involving the Decision Token and training on synthetic non-function-call data are discussed in Section 4.3.
3.3 Chain-of-thought Reasoning
CoT reasoning has been demonstrated to significantly enhance performance across various tasks by incorporating intermediate reasoning steps Wei et al. (2022). Inspired by this, we explore whether CoT reasoning can similarly improve function-calling capabilities. To achieve this, we propose a synthetic data generation pipeline that constructs reasoning descriptions derived from sequences of conversations and function calls. This pipeline leverages single-turn queries with commercial-grade LLMs. In our prompt design, we initially provide the history of the conversation and the available functions, requiring the identification of the reasoning needed to determine how to use the available functions to achieve the target function calls. Additionally, we provide multiple examples to enhance stability (few-shot learning). More details are provided in Appendix A. Using this pipeline, we generate data that captures the thinking process, which is then used to fine-tune base LLMs. The fine-tuning process employs a structured prompt template, as illustrated in Figure 1(h). The experiments on incorporating CoT reasoning are presented in Section 4.4.
3.4 Multilingual Translation
To enhance the multilingual capabilities of function-calling tuning, translating existing English function-calling datasets into target languages is a common approach. However, this process presents significant challenges, as elements such as function names, enumeration items, and structured function calls cannot be directly translated without risking inconsistencies or errors. To address these issues and maintain the semantic and syntactic integrity of translated datasets, we propose a novel translation pipeline specifically designed to overcome the limitations of direct translation methods. This pipeline leverages a single-turn query with commercial-grade LLMs. In our prompt design, we initially specify the JSON format of the provided conversation trajectory with function calls. We then instruct the LLMs to translate the data into the target language, adhering to the rules of not translating function names and descriptions, and translating arguments only when reasonable. More details are provided in Appendix B. The experiments on verifying the effectiveness of the pipeline is presented in Section 4.5.
Use of Data? MT AST Relevance
Conditional Prompt IF-110k FC-110k Bench Summary Detection
[…]
Table 1: Performance comparison of different prompts and the use of data on various metrics for instruction-following and function-calling capabilities. The "Use of Data?" columns indicate whether the respective datasets (IF-110k and FC-110k) are included in the training process. Detailed experiments are discussed in Section 4.2.
How to provide functions in a prompt? In a dedicated role In the system role
Metrics on BFCL Yan et al. (2024): AST Relevance AST Relevance
Summary Detection Summary Detection
Baseline 85.25 49.58 85.94 39.58
+ Decision Token 85.25 37.50 84.63 47.50
+ Non-function-call Data (NF-1k) 84.81 57.50 83.44 65.42
Table 2: Impact of incrementally adding the Decision Token and synthetic non-function-call data. The table shows different prompt configurations for providing functions. The last three rows represent the configurations: baseline, Decision Token added, and both Decision Token and synthetic data added. See Section 4.3 for details.
4 Experiments and Results
4.1 Experimental Setup
In this section, we describe the experimental setup used to evaluate our proposed methods, including details on datasets, model configurations, training parameters, and evaluation metrics.
We created a diverse dataset for fine-tuning, which includes both instruction-following and function-calling examples. The instruction-following data, marked as IF-110k, consists of 110k instances sampled from Open ORCA Longpre et al. (2023), a synthetic dataset generated from GPT-4 completions. The function-calling data, marked as FC-110k, also includes 110k instances, sourced from a combination of APIGen Liu et al. (2024b) and the glaive-function-calling-v2 dataset [^1] https://huggingface.co/datasets/glaiveai/glaive-function-calling-v2.
We used Breeze-7B [^2] https://huggingface.co/MediaTek-Research/Breeze-7B-Base-v1_0 as the base model for our experiments. Breeze-7B Hsu et al. (2024) is an open-source language model based on Mistral-7B, designed to improve language comprehension and chatbot capabilities in Traditional Chinese. Using Breeze-7B, we can test the model’s effectiveness in both English and Traditional Chinese.
The models were fine-tuned using the prompt templates, described in Section 3.1. For fine-tuning, we applied the low-rank adaptation (LoRA) technique on linear layers. The fine-tuning process used the following hyperparameters: a learning rate of 1e-4, a batch size of 48, 3 epochs, a cosine learning rate scheduler, the AdamW optimizer, 100 warmup steps, a LoRA rank ($r$) of 16, and a LoRA $\alpha$ of 32.
We evaluated the performance of our models using the following metrics:
AST Summary (%): This metric, used in the Berkeley Function Calling Leaderboard (BFCL) Yan et al. (2024), assesses the structural correctness of language model outputs for function-calling tasks by comparing the Abstract Syntax Tree (AST) representations of generated and target function calls. It includes four problem types—Simple Function, Multiple Function, Parallel Function, and Parallel Multiple Function—categorized based on the combination of the number of provided functions and function calls. The dataset consists of 400 Simple Function tasks and 200 tasks for each of the other three types. The AST Summary is the average accuracy across these four types.
Relevance Detection (%): This metric, also used in the BFCL, measures the success rate of no function call when none of the provided functions are relevant. This scenario helps determine whether a model will hallucinate its functions and parameters when the provided functions are irrelevant to the user’s query.
MT-Bench (score): Unlike previous works, we also explore the impact of instruction-following capabilities when enabling function-calling functionalities. MT-Bench Zheng et al. (2023) is a benchmark for evaluating these capabilities. We use GPT-4o as a judge to give the score out of 10.
We also evaluated the performance on Traditional Chinese function calling using the Function Calling Leaderboard for ZHTW Lee et al. (2024), which is constructed by translating the BFCL. Therefore, the calculation of metrics AST Summary and Relevance Detection is similar.
How to provide functions in a prompt? In a dedicated role In the system role
Metrics on Function Calling Leaderboard AST Relevance AST Relevance
   for ZHTW Lee et al. (2024): Summary Detection Summary Detection
Baseline 52.37 36.67 50.81 47.08
+ Traditional Chinese Data (TC-19k) 61.56 41.25 58.56 45.83
Table 3: The impact of adding Traditional Chinese data, generated through a tailored translation pipeline (Section 3.4), is analyzed. Notably, the metrics AST Summary and Relevance Detection are evaluated on the benchmark for Tradition Chinese. Detailed experiments are discussed in Section 4.5.
4.2 Effects of Prompt Templates and Use of Training Data
We investigated the performance comparison of different conditional prompts and the use of training data on various metrics for instruction-following and function-calling capabilities, as shown in Table 1. Conditional prompts are described in Section 3.1. The use of training data, training setup, and metrics is described in Section 4.1.
Compared to Table 1(b) and (c), the functions provided in a dedicated role and the system role exhibit similar capabilities in terms of instruction-following (MT Bench) and function-calling accuracy (AST Summary). But, Relevance Detection is superior when functions are provided in the dedicated role. We hypothesize that providing functions in the dedicated role makes the template with functions significantly different from the template without functions, making it easier for the model to learn when to use function calling or respond directly.
Compared to the results shown in Table 1(a), (b), and (c) on the MT Bench, we find that enabling the function-calling capability does not reduce the performance of the instruction-following capability, regardless of the conditional prompt given.
Compared to the results shown in Table 1(b), (c), (d), and (e) on the AST Summary and Relevance Detection metrics, we find that the performance of the function-calling capability decreases when we exclude the instruction-following data (IF-110k). This observation is noteworthy. We hypothesize that the increase in function-calling capability is due to the additional instruction-following data, which helps the model better understand the semantic structure of the prompts. Consequently, this improved understanding enhances the model’s ability to accurately perform function calling. Moreover, instruction-following data provided more non-function-call examples, further improving Relevance Detection.
In conclusion, our experiments demonstrate that the inclusion of function-calling capabilities does not compromise instruction-following performance. Additionally, the use of instruction-following data significantly enhances function-calling accuracy and relevance detection.
4.3 Effects of the Decision Token
To verify the effectiveness of the Decision Token, as described in Section 3.2, we examined the effects of incrementally adding the Decision Token and the synthetic non-function-call data.
In the baseline experiment, we used IF-110k and FC-110k as the training data to finetune the base model. Then, we added the Decision Token to the prompt templates and finetuned the base model on the same training data. In the final experiment, we used synthetic methods described in Section 3.2 to generate 1k instances of non-function-call data, marked as NF-1k. The models were then finetuned with a combination of NF-1k, IF-110k, and FC-110k. The results of this investigation are presented in Table 2. In conclusion, our analysis shows that the adoption of the Decision Token, along with the accompanying synthetic non-function-call data, can benefit Relevance Detection. However, it also results in a slight decrease in function-calling accuracy (AST Summary).
4.4 Effects of Chain-of-Thought Reasoning
To evaluate CoT reasoning (Section 3.3), we generated reasoning descriptions for each function call in FC-110k, creating FC-110k-Reason. Comparing models trained on IF-110k + FC-110k-Reason with those trained on IF-110k + FC-110k, we found no significant improvement in function calling accuracy (AST Summary), which was 84.44% compared to the baseline of 85.25%. We hypothesize that BFCL problems may not require reasoning for function calling.
4.5 Effects of Translation Pipeline
To evaluate the effectiveness of the translation pipeline described in Section 3.4, we generated 18k function-calling instances in Traditional Chinese using synthetic methods from the FC-110k dataset. Additionally, we applied an non-function-call case generation pipeline, as detailed in Section 3.2, to this dataset, producing 200 instances of non-function-call data in Traditional Chinese. The combined dataset is referred to as TC-19k.
In our baseline experiment, we used the Decision Token approach along with the IF-110k, FC-110k, and NF-1k datasets as training data. We then incorporated the TC-19k Traditional Chinese data into the training set. The results, presented in Table 3, demonstrate that even a small amount of translated data can significantly enhance function-calling performance.
5 Conclusion
Our research demonstrates that integrating instruction-following data with function-calling tasks significantly enhances function-calling capabilities. The Decision Token mechanism, combined with synthetic non-function-call data, further improves relevance detection. Additionally, a tailored translation pipeline effectively mitigates multilingual challenges. These findings underscore the potential for improving function-calling capabilities and expanding multilingual proficiency in LLMs, paving the way for more practical real-world applications.
Appendix A Details of pipeline for constructing reasoning descriptions
The following prompt is for constructing reasoning descriptions. The provided conversation trajectory is given in {CONVERSATIONS}, the provided function descriptions are in {FUNCTIONS}, and the provided function calls are in {FUNC_CALL}.
     Your mission is to identify the reason for using the tool based on the history conversations.    ## Example 1:    Given the history conversations as follows:  """  [SYSTEM] You are a helpful assistant.  [USER] What is the weather in Taipei?  [BOT] Current temperature in Taipei: 32 Celsius  [USER] What is the weather in Palo Alto?  """  and the available tools are as follows:  ```json  [    {      "name": "weather_api.get_current_weather",      "description": "Retrieves the current weather conditions for a specified location.",      "parameters": {        "location": {          "type": "string",          "description": "The name of the city or geographic location.",          "required": true        },        "units": {          "type": "string",          "description": "The units for temperature measurement (e.g., ’Celsius’, ’Fahrenheit’).",          "required": false        }      }    }  ]  ```    Please output JSON with the key `reason` for identifying the reason  to figure out how to use the available functions and finally expect to get the answer shown below.  ```json  [    {      "name": "weather_api.get_current_weather",      "arguments": {        "location": "Palo Alto",        "units": "Celsius"      }    }  ]  ```    ## Output for Example1    ```json  {    "reason": "The user wants to know the current weather conditions in Palo Alto. The available tool ’weather_api.get_current_weather’ can be used to retrieve this information by specifying the location as ’Palo Alto’."  }  ```    ## Example 2:    Given the history conversations as follows:  """  [USER] Find the sum of all the multiples of 3 and 5 between 1 and 1000. Also find the product of the first five prime numbers.  """  and the available tools are as follows:  ```json  [    {      "name": "math_toolkit.sum_of_multiples",      "description": "Find the sum of all multiples of specified numbers within a specified range.",      "parameters": {        "lower_limit": {          "type": "integer",          "description": "The start of the range (inclusive).",          "required": true        },        "upper_limit": {          "type": "integer",          "description": "The end of the range (inclusive).",          "required": true        },        "multiples": {          "type": "array",          "description": "The numbers to find multiples of.",          "required": true        }      }    },    {      "name": "math_toolkit.product_of_primes",      "description": "Find the product of the first n prime numbers.",      "parameters": {        "count": {          "type": "integer",          "description": "The number of prime numbers to multiply together.",          "required": true        }      }    }  ]  ```    Please output JSON with the key `reason` for identifying the reason  to figure out how to use the available functions and finally expect to get the answer shown below.  ```json  [    {      "name": "math_toolkit.sum_of_multiples",      "arguments": {        "lower_limit": 1,        "upper_limit": 1000,        "multiples": [3, 5]      }    },    {      "name": "math_toolkit.product_of_primes",      "arguments": {        "count": 5      }    }  ]  ```    ## Output for Example2    ```json  {    "reason": "The user wants to find the sum of all multiples of 3 and 5 between 1 and 1000, and also find the product of the first five prime numbers. The available tools ’math_toolkit.sum_of_multiples’ and ’math_toolkit.product_of_primes’ can be used to retrieve this information. The ’math_toolkit.sum_of_multiples’ tool can be used by specifying the lower limit as 1, the upper limit as 1000, and the multiples as [3, 5]. The ’math_toolkit.product_of_primes’ tool can be used by specifying the count as 5."  }  ```    ## Start    Given the history conversations as follows:  """  {CONVERSATIONS}  """  and the available tools are as follows:  ```json  {FUNCTIONS}  ```    Please output JSON with the key `reason` for identifying the reason  to figure out how to use the available functions and finally expect to get the answer shown below.  ```json  {FUNC_CALL}  ```
Appendix B Details of pipeline for translating function-calling data
The following prompt is for translating function-calling data, where the provided function-calling data in JSON format is specified in {DATA}, and the target language is indicated in {TARGET_LANG}, e.g., "Traditional Chinese."
     This JSON object outlines a conversation between a user and an assistant, including the available functions the assistant can utilize to meet the user’s requests.    In this JSON object:  - The `functions` key lists the available functions the assistant can use, including their descriptions and parameters.  - The `conversations` key outlines the conversation between the user and the assistant.  - The `tool_calls` key within the assistant’s response shows the function calls the assistant makes to fulfill the user’s requests, including the function name and arguments.    ```json  {DATA}  ```    AND NOW,  I want to translate this JSON into {TARGET_LANG}.  Note that:  - Do not translate any content in `functions`  - Translate the content in `arguments` if using {TARGET_LANG} is reasonable    Please provide your translation into JSON as same format above.